OpenStudy (anonymous):

Solve using the quadratic formula. x² + x = 12

OpenStudy (anonymous):

\[x=-b+/-\sqrt(b^2-4ac)/2a\]

OpenStudy (anonymous):

plug in a=1, b=1, c=-12

OpenStudy (eyust707):

\[x = {-b \pm \sqrt{b^2 - 4ac} \over 2a}\]

OpenStudy (anonymous):

^how did you do that :O

OpenStudy (anonymous):

thx

OpenStudy (eyust707):

In the equation editor i wrote: x = {-b pm sqrt{b^2 - 4ac} over 2a}

OpenStudy (eyust707):

the out side most brackets are for the "over" so that they dont mess with the "x ="

OpenStudy (anonymous):

:O Thank you! I'll remember that.

OpenStudy (eyust707):

=P heres another trick... If you want to include symbols like \(\pi\) or \(\sqrt{112}\) within your text without starting a new line you can do so by replacing the [ ] with ( )

OpenStudy (anonymous):

Thanks!